In a large mixing bowl, sugar is poured in at a rate of 2t + 5 cubic centimeters per
second, and milk is mixed in at a rate of 2t

2 + 1 cubic centimeters per second. After
how long does the amount of milk exceed the amount of sugar in the bowl?
1
Expert's answer
2020-04-24T20:09:24-0400

Let "S(t)=2t+5" - amount of sugar

and "M(t)=2t^2+1" - amout of milk

We should solve following inequality:

"M(t)>S(t)"

"2t^2+1>2t+5 \\ 2t^2-2t-4>0\\ t^2-t-2>0"

"D=(-1)^2-4(-2)=9"

"t_1=\frac{1+3}{2}=2\\
t_2=\frac{1-3}{2}=-1"


"t\isin(-\infin,-1)\bigcup(2,+\infin)"

Since "t" is time, it can't be less than 0.

Thereby, amount of time will exceed amount of sugar when "t>2"



Here blue line shows amount of sugar and red line shows amount of milk.